OpenStudy (anonymous):

I need the answer so bad |8x + 10| = 90

hero (hero):

8x + 10 = 90 8x + 10 = -90

OpenStudy (mertsj):

That sentence means that 8x+10 is exactly 90 units from 0

OpenStudy (mertsj):

Therefore it must be 90 or -90

OpenStudy (mertsj):

So write that 8x+10=90 or 8x+10=-90 Solve each equation to find that x = 80/8=10 or 8x=-100/8=-25/2

OpenStudy (mertsj):

so the two answers are 10 and -25/2

OpenStudy (anonymous):

My mistake there, well done MertsJ!